OpenStudy (peterg2001):

Which of the following sets of sides does not form a right triangle? 6, 8, 10 8, 15, 17 7, 24, 26 5, 12, 13 @Molly_Qian

OpenStudy (anonymous):

so 24 26 isn't

OpenStudy (anonymous):

the one that contains *

OpenStudy (peterg2001):

???/??!@Molly_Qian

OpenStudy (peterg2001):

@EclipsedStar @kirbykirby

OpenStudy (anonymous):

@peterg2001, what condition can you check to prove something is a right angled triangle?

OpenStudy (peterg2001):

A N S W E R ?????????????

OpenStudy (solomonzelman):

Whether a set of site belongs to a right triangle or not can be checked in a failry easy way. For a set of sides \(\large\color{black}{ \displaystyle {\rm \color{red}{a}},~~{\rm \color{green}{b}},~~{\rm \color{blue}{c}} }\) \(\large\color{black}{ \displaystyle ( }\)where \(\large\color{black}{ \displaystyle{\rm \color{blue}{c}} }\) is the biggest side\(\large\color{black}{ \displaystyle) }\) the following has to be true: \(\large\color{black}{ \displaystyle {\rm \color{red}{a}}^2+{\rm \color{green}{b}}^2={\rm \color{blue}{c}}^2 }\)

OpenStudy (anonymous):

@peterg2001 the one that contains 7, 24, 26 is not a right trangle